OpenStudy (anonymous):

A baseball of mass 100 grams is traveling 30m/s when it is caught. In catching the ball player moves his mitt backwards 21 cm and stops. What average force does the ball exert when it is caught?

OpenStudy (michele_laino):

here, the work done by the force applied by the player, has to be equal to the kinetic energy of the ball. So we can write: \[\frac{1}{2}m{v^2} = Fd\] where \(d=21\)cm

OpenStudy (michele_laino):

from the formula above, i get: \[F = \frac{{m{v^2}}}{{2d}} = \frac{{0.1 \cdot {{30}^2}}}{{2 \cdot 0.21}} = ...?\]

OpenStudy (anonymous):

F=214.285Newton's